# Primordial Power Spectrum of Five Dimensional Uniform Inflation

### v2

@@ Abstract
Five dimensional (5D) uniform inflation describes a de Sitter (or approximate) solution of 5D Einstein equations, with cosmological constant and a 5D Planck scale $M_{*}\sim 10^{9}~{}{\rm GeV}$. During the inflationary period all dimensions (compact and non-compact) expand exponentially in terms of the 5D proper time. This set-up requires about 40 $e$-folds to expand the fifth dimension from the fundamental length to the micron size. At the end of 5D inflation (or at any given moment during the inflationary phase) one can interpret the solution in terms of 4D fields using 4D Planck units from the relation $M_{p}^{2}=2\pi RM_{*}^{3}$, which amounts going to the 4D Einstein frame. This implies that if the compactification length $R$ expands $N$ $e$-folds, then the 3D space would expand $23N/$ 2 $e$-folds as a result of a uniform 5D inflation. We reexamine the primordial power spectrum predicted by this model and show that it is consistent with Planck’s measurements of the comic microwave background. The best-fit to Planck data corresponds to $R\sim 10~{}\mu{\rm m}$. A departure of the angular power spectrum predicted by 4D cosmology is expected at multipole moment $\ell\sim 7$.
@@
where $A_{s}\simeq 2\times 10^{-9}$ is the scalar amplitude, $n_{s}\simeq 0.96$ is the scalar spectral tilt, and $k_{*}=0.05~{}{\rm Mpc}^{-1}$ is the pivot scale that exits the horizon at $N_{*}\equiv N_{\rm end}-50$ e-folds from the start of inflation ($N_{\rm start}\equiv 0$Planck:2018nkj . [^1] A point worth noting at this juncture is that any given present-day value “$x$” of a CMB scale $k$, including the pivot scale $k_{*}$, should be expressed in the form of $k=a_{\rm today}\,x~{}{\rm Mpc}^{-1}$. However, it is generally assumed that $a_{\rm today}=1$ implicitly, and the scale is expressed as $k=x~{}{\rm Mpc}^{-1}$ Mishra:2021wkm .
@@
$\Omega_{b}h^{2}$ [0.02, 0.0265]0.020.0265[0.02,\,0.0265][ 0.02 , 0.0265 ]
$\Omega_{c}h^{2}$ [0.1, 0.135]0.10.135[0.1,\,0.135][ 0.1 , 0.135 ]
$100\theta_{s}$ [1.03, 1.05]1.031.05[1.03,\,1.05][ 1.03 , 1.05 ]
$\tau_{\rm reio}$ [0.03, 0.08]0.030.08[0.03,\,0.08][ 0.03 , 0.08 ]
$n_{s}$ [0.920, 0.996]0.9200.996[0.920,\,0.996][ 0.920 , 0.996 ]
$\ln(10^{10}A_{s})$ [2.763, 4.375]2.7634.375[2.763,\,4.375][ 2.763 , 4.375 ]
$\ln(10^{10}{\cal A}_{s})$ [2.763, 4.375]2.7634.375[2.763,\,4.375][ 2.763 , 4.375 ]
$\zeta$ $[-6.6,\,-2.1]$
@@
In Fig. 1 we show a comparison between the primordial power spectra of the concordance 4D model of cosmology (dashed line) and 5D uniform inflation (solid line) from the likelihood analysis of Ref. Petretti:2024mjy . One can check by inspection that the primordial power spectrum predicted by 5D uniform inflation departs from the 4D $\Lambda$CDM prediction at around $k\sim 2/{\cal R}_{\rm today}\sim 5\times 10^{-4}~{}{\rm Mpc}^{-1}$. This corresponds to a multipole moment $\ell\simeq kd_{A}\simeq 7$, where $d_{A}=1.4\times 10^{4}~{}{\rm Mpc}$ is the angular diameter distance to the last scattering surface Bridle:2003sa . An explicit comparison between the different interpretations of the result from the likelihood analysis is provided in Appendix II.
@@
Note Added
Within the dark dimension scenario cosmological overproduction of bulk graviton modes leads to a normalcy temperature $T_{*}\sim 1~{}{\rm GeV}$ Arkani-Hamed:1998sfv ; Gonzalo:2022jac . The behavior of the scale factor between the end of inflation and $T_{*}$ is model dependent. In (16) we adopted a simple radiation-like behavior for the scale factor between $M_{I}$ and $T_{*}$, but of course assuming that the universe must be free of bulk modes above $T_{*}$. As noted in Macesanu:2004gf , one possible pathway to accommodate our assumption could be the softening of the gravity-matter interaction due to brane fluctuations Bando:1999di ; Bando:2000ch . Indeed, if brane oscillations were to be excited above the normalcy temperature, then the matter-graviton interaction vertex would aquire an effective form factor ${\cal F}=e^{-\frac{1}{2}\frac{m^{2}}{\Delta^{2}}}$, where $m$ is the mass of the Kaluza-Klein graviton and $\Delta\sim 4\pi\sqrt{\tau}/M_{*}$ is the “softening scale”, which is related to the brane tension $\tau$ Murayama:2001av . As a consequence, the cross-section for production of gravitons could be exponentially suppressed with a suitably small tension of our world, i.e., for a relatively soft brane. It is interesting to note that the cooling process of Supernova 1987A leads to the constraint $\tau\gtrsim 100~{}{\rm GeV}^{4}$ Kugo:1999mf . An investigation along this line is obviously important to be done.

### v3

@@ Abstract
Five dimensional (5D) uniform inflation describes a de Sitter (or approximate) solution of 5D Einstein equations, with cosmological constant and a 5D Planck scale $M_{*}\sim 10^{9}~{\rm GeV}$. During the inflationary period all dimensions (compact and non-compact) expand exponentially in terms of the 5D proper time. This set-up requires about 40 $e$-folds to expand the fifth dimension from the fundamental length to the micron size. At the end of 5D inflation (or at any given moment during the inflationary phase) one can interpret the solution in terms of 4D fields using 4D Planck units from the relation $M_{p}^{2}=2\pi RM_{*}^{3}$, which amounts going to the 4D Einstein frame. This implies that if the compactification length $R$ expands NN $e$-folds, then the 3D space would expand $3N/2$ $e$-folds as a result of a uniform 5D inflation. We reexamine the primordial power spectrum predicted by this model and show that it is consistent with Planck’s measurements of the comic microwave background. The best-fit to Planck data corresponds to $R\sim 10~\mu{\rm m}$. A departure of the angular power spectrum predicted by 4D cosmology is expected at multipole moment $\ell\sim 7$.
@@
where $A_{s}\simeq 2\times 10^{-9}$ is the scalar amplitude, $n_{s}\simeq 0.96$ is the scalar spectral tilt, and $k_{*}=0.05~{\rm Mpc}^{-1}$ is the pivot scale that exits the horizon at $N_{*}\equiv N_{\rm end}-50$ e-folds from the start of inflation ($N_{\rm start}\equiv 0$Planck:2018nkj . [^1] A point worth noting at this juncture is that any given present-day value “xx” of a CMB scale $k$, including the pivot scale $k_{*}$, should be expressed in the form of $k=a_{\rm today}\,x~{\rm Mpc}^{-1}$. However, it is generally assumed that $a_{\rm today}=1$ implicitly, and the scale is expressed as $k=x~{\rm Mpc}^{-1}$ Mishra:2021wkm .
@@
$\Omega_{b}h^{2}$ [0.02, 0.0265][0.02,\,0.0265]
$\Omega_{c}h^{2}$ [0.1, 0.135][0.1,\,0.135]
$100\theta_{s}$ [1.03, 1.05][1.03,\,1.05]
$\tau_{\rm reio}$ [0.03, 0.08][0.03,\,0.08]
$n_{s}$ [0.920, 0.996][0.920,\,0.996]
$\ln(10^{10}A_{s})$ [2.763, 4.375][2.763,\,4.375]
$\ln(10^{10}{\cal A}_{s})$ [2.763, 4.375][2.763,\,4.375]
$\zeta$ [6.6,2.1][-6.6,\,-2.1]
@@
[…]